Two camper vans leave Arrowhead Lake at the same time, one traveling north at a speed of 10 km/h faster than the other, which is traveling south. After 3 hours, the camper vans are 420 km apart. Find their speeds.
Answer by stanbon(75887)

Two camper vans leave Arrowhead Lake at the same time, one traveling north at a speed of 10 km/h faster than the other, which is traveling south. After 3 hours, the camper vans are 420 km apart. Find their speeds.
---------------
South van DATA:
rate = x kph; time= 3 hrs.; distance = 3x km
------------
North van DATA:
rate = x+10 kph; time = 3 hrs; distance = 3(x+10)=3x+30 km
---------------
EQUATION:
dist. + dist. = 420 km
3x+3x+30=420
6x=390
x= 65 kph (speed of south-bound van)
x+10= 75 (speed of north-bound van)
Cheers,
Stan H.